A) Back Alley Boys, Inc. had sales of $250,000, cost of goods sold of $80,000, depreciation expense of $27,000, and additions to retained earnings of $33,360. The firm paid out $30,000 in dividends. Assume a 34% income tax rate, what is the times interest earned ratio?



B) Using the following data, compute the value of cost of goods sold for Peterson Brewing:


Current liabilities=$340,000 quick ratio=1.8 inventory turnover=4.0 current ratio=3.3



1
Expert's answer
2021-11-21T17:35:19-0500

A)

Times interest earned ratio can be calculated by dividing a company’s EBIT by its periodic interest expense.

TIE Ratio = (Earning Before Interest and Taxes)/(Interest Expense)

Gross profit = Sales - COGS

=250,000 - 80000

=$170000

Thus, EBIT= Gross Profit - Depreciation Expenses + Retained Earnings - Dividents

= 170000 - 27000 + 33,360 - 30000

= $146360

Interest Expense = Total Debt * Interest rate

= 30000*0.34

= $10200

Hence, TIE Ratio = 146360/10200

=14.3x

B)

Given;

Current liabilities = $340,000

Quick ratio = 1.8

Inventory turnover = 4.0

Current ratio = 3.3.

COGS=Beginning Inventory+P−Ending Inventory

where

P=Purchases during the period

​ =(340000/1.8) + 340000 -(340000/4)

= 188888 + 340000 - 85000

COGS=$443888